OpenStudy (anonymous):

What's 5 and 1/7 ÷ 9?

OpenStudy (anonymous):

(1/7 ) / 9= 1/63

OpenStudy (anonymous):

5 / (9)= 9/5

OpenStudy (anonymous):

-2660921/-4656612